# Self-Consistent Stochastic Finite-Temperature Modelling: Ultracold Bose Gases with Local (s-wave) and Long-Range (Dipolar) Interactions

### v1

@@ IV.3.1 Stochastic Approach of Stoof
Before commenting further on the use of such an equation in Sec. IV.3.3, we note here that an equation like (44) was also obtained as the limiting case of an alternative approach, briefly discussed below.
IV.3.2 Stochastic Approach of Gardiner et al.
@@
[…]
@@
In the SPGPE formalism, the contribution associated with $G(r)$ is known as a growth term (consistent with Stoof’s earlier discussion), and corresponds to an energy- and momentum-conserving binary collision which leads to a change in the number of particles within the $\Phi_{0}$ band. The second term, associated with $M$, corresponds to a collision between a coherent band and an incoherent band particle, which shifts the energies, but does not lead to net population transfer between the two bands. As such, these terms – whose explicit expressions can be found in [20] – have been respectively referred to as ‘number-damping’ and ‘energy-damping’ contributions [101, 102, 103, 104].
@@
it is easy to see that both formalisms are equivalent, upon also dropping the explicit appearance of the projector of the SPGPE formalism. Note that as both ourselves and Stoof are working in a path integral formalism, such projectors do not appear explicitly, but the existence of these two distinct subspaces is implicit within the path integral approach, becoming relevant in numerical implementations.
Ignoring the so-called scattering term reproduces Eq. (44) (but now with the projector, $\mathcal{P}$, appearing explicitly in the equation), and after redefining $\psi\to\psi e^{i\mu t}$ and $dW_{G}\to dW_{G}\,e^{i\mu t}$, thus leading to the so-called ‘simple growth’ SPGPE, having the form
@@
IV.3.3 Stochastic (Projected) Gross-Pitaevskii Predictions
@@
IV.3.4 The Dissipative (or Phenomenologically Damped) Gross-Pitaevskii Equation
@@
V Corresponding Finite-Temperature Equations for Dipolar Gases
In the previous section, we have explicitly shown how our generalised stochastic formalism provides an appealing, more transparent, unifying way, to obtain a broad range of established theories for weakly-interacting atomic gases with contact interactions as limiting cases, based on different approximations. Most notably, our formalism can be seen as creating a unifying framework simultaneously incorporating both stochastic growth and scattering processes on the one hand, and quantum Boltzmann thermal particle dynamics on the other hand, in a dynamically self-consistent manner.
In this section, however, we demonstrate the full power of our methodology, by extending our model to systems interacting with long-range interactions: specifically, we consider a typical example from the ultracold atomic gas community in the form of dipolar condensates, whose typical interatomic potential contains, in addition to the local ($\delta$-function) contributions also a long-range interaction of the form of Eq. (3). As mentioned in the introduction, our parallel handling of gravitational attraction treated as a Poisson equation [107], has shown us that a long-range interaction – such as the dipolar one – should also in principle be amenable to modelling via an additional Poisson-like equation for the dipolar potential, as explained below.
V.1 The Gross-Pitaevskii Equation
@@
This equation describes the mean-field regime of dipolar gases in the near-$T=0$ limit. Recently, significant attention has been placed in modelling so-called dipolar droplets [179, 180], i.e. self-bound ‘mini-condensates’ whose collapse is stabilised by quantum fluctuations through a quantum correction term to the mean-field energy, known as the Lee-Huang-Yang, or LHY, contribution [58, 169, 170, 171, 172]. Such beyond-mean-field physics is clearly not contained within the above GPE. Nonetheless, there exists a generalized version of it, known as the extended GPE, or eGPE, which has quickly become the norm in this community [169, 170, 171, 172, 175, 176, 162], and which we therefore briefly discuss below.
V.1.1 The Extended Gross-Pitaevskii Equation (eGPE) for Dipolar Condensates
Following the work of Lima and Pelster [169, 170]) (see Appendix C for details), one can linearize the condensate wavefunction, following the usual Bogoliubov-de Gennes analysis, via
@@ V.1.1 The Extended Gross-Pitaevskii Equation (eGPE) for Dipolar Condensates
This section has outlined an established proof of how the equation for the pure condensate mean-field can be systematically extended to account for the role of quantum fluctuations, in the regime where such terms become important. However, it is pertinent to remind the reader that all other formal considerations within this work have focussed on the rather distinct regime where thermal fluctuations dominate. It is therefore plausible to expect that such two distinct pictures, valid in separate limits of the relative ratio of importance of thermal to quantum effects, can be somehow combined, and we comment on this further below (Sec. V.3).
However, before doing so, we should further clarify why such quantum correction terms do not emerge naturally from our presented formalism: In this context we note that our Keldysh non-equilibrium approach has focussed on including the effects of fluctuations of the fast (non-coherent) fields to all orders, but has only considered stochastic/statistical fluctuations of the slow fields, arising from the influence of the fast fields. Furthermore, our analysis so far was done within the so-called ‘Popov approximation’ [56]: such an approximation amounts to ignoring anomalous average contributions within our approach. Such anomalous contributions are well-known to lead to further many-body effects and quantum depletion (see also, e.g., [62, 64]), typically expressed in terms of the parameter $(na^{3})$ which characterizes their relative importance. In typical alkali gases, $(na^{3})\ll 1$, and such terms can, in most cases, be ignored to good approximation. While the same is expected to be true for dipolar gases in the regime of dominant thermal fluctuations, there are interesting, experimentally-accessible regimes where this no longer holds, with such quantum corrections dominating, for example, the physics of dipolar droplets and supersolids. For a discussion of the dipolar gas phase diagram, and such distinct regimes, we refer the reader to a recent review [162].
While such extra quantum-correction terms can in fact enter our formalism via the additional incrorporation of quasi-particle physics (see Appendix C for further details), such handling is additional to our presented Keldysh path integral formulation based on a quartic Hamiltonian. We conclude this discussion by noting in passing that, while our selected approach does not seem capable to generate such quantum correction terms without such additional action in the standard limit of the binary interaction Hamiltonian of Eq. (1), such terms would be expected to arise, at least in principle, when starting from a more general Hamiltonian including more than two particle effects [168] – as might be expected to be relevant in denser samples beyond the dilute gas limit.
V.2 Gross-Pitaevskii-Boltzmann, or ZNG Model of Dipolar Condensates
@@
We stress that the equations presented in this section arise directly from a formalism explicitly focussing on thermal fluctuations: in the regime where such thermal fluctuations are important, leading to corrections on the otherwise pure-condensate physics, it might be reasonable to expect that quantum fluctuations will typically be sub-dominant. As such, it might be that the most relevant regimes in ultracold dipolar systems are either those with condensate and thermal fluctuations (such as the dipolar ZNG equations of Eqs. (70)-(72)), or those with condensate and quantum fluctuations (such as the extended GPE of Eq. (67)), essentially choosing between the generalised finite-temperature models self-consistently derived in an ab initio manner in our work, or, at lower temperatures, to the established LHY quantum correction term summarized in Sec. V.1.1 and leading to the extended GPE of Eq. (67).
@@
V.3 An Extended Gross-Pitaevskii, Collisionless Boltzmann Equation Model?
Considering the somewhat ‘orthogonal’ directions of travel presented in Sec. V.1.1 (quantum fluctuations) and Sec. V.2 (thermal fluctuations) – each expected to be valid in the relevant regime –, and noting that the latter approach on which our main equations are based does not include condensate fluctuations to all orders, one might consider postulating a combination of those two pictures, albeit in a somewhat ad hoc manner, to interpolate between such regimes.
@@ V.3 An Extended Gross-Pitaevskii, Collisionless Boltzmann Equation Model?
Such a parameter separates the regimes of sub-dominant quantum fluctuations (${\cal F}\ll 1$) in which our ab initio formalism emerges naturally, from that when they dominate (${\cal F}\gg 1$). We note here that while the net effect will be given by $\int d^{3}{\bf r}\,{\cal F({\bf r})}$, the inhomogeneous densities (droplets) formed by the LHY term may in fact imply that its effect cannot be locally ignored even if $\int d^{3}{\bf r}{\cal F({\bf r})}\ll 1$ overall in the gas. As well-known the parameter characterizing their relative importance is $g-C_{dd}/3=g(1-\epsilon_{dd})$. As $\epsilon_{dd}\rightarrow 1$, as is the case in various dipolar condensate regimes, ${\cal F}({\bf r})\gg 1$, and so the above postulated contribution $+\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}$ becomes highly-relevant, even if only locally so.
Alternatively, observation of Eq. (76) shows that, for $C_{dd}/3\ll g$, or equivalently $\epsilon_{dd}\ll 1$, and a reasonable intermediate temperature $0<T<T_{c}$, the quantum fluctuation term may only pose a minor correction to the overall system properties and dynamics.
The interesting and pertinent question of course remains, whether the above postulated combination of quantum and thermal contributions [Eqs. (73)-(75)], which is known to work well in the respective quantum-dominated and thermal-dominated limits may actually provide valid predictions even in the intermediate regime when ${\cal F}\sim O(1)$. In the absence of any other advanced theoretical models addressing this, it may thus be relevant to investigate that numerically, which we defer to a future publication.
V.4 (Simple Growth) Stochastic Gross-Pitaevskii Equation for Dipolar Condensates
@@
As discussed earlier [Sec. IV] for the weakly-interacting gas limit, such an equation is not useful to work with on its own, as the term $iR$ is dynamical, encompassing the dynamics of the incoherent modes.
@@
where the dissipation parameter $\gamma$ is defined here in the usual manner as
@@
One important point that emerges from this equation relates to the explicit appearance of the non-condensate density contribution $\tilde{n}$ in Eqs. (82)-(82). Previously, in the contact-interaction setting, we argued that while any quantitative comparison to experiments strictly requires that contribution to be included, ignoring , to lowest approximation, such a contribution might still give reasonable (but perhaps only qualitative) results, with such treatment possibly also able to model experiments (semi)quantitatively when further supplemented with an appropriately fit $\gamma$ parameter, obtained from experimental measurements. Contrary to that, in the dipolar case we feel it is pertinent to highlight that the situation is more complicated. The reason for this is that the effective dipolar potential $V_{d}$, included in its integrated form in Eq. (82), actually depends critically on the thermal density itself, due to the long-range nature of the interactions. As such, it is not a priori obvious that there exist interesting finite-temperature regimes where ignoring such contributions can be valid. A natural question of course arises how such non-condensate densities could in fact be included in the treatment. Borrowing ideas from earlier treatments, it might be valid in some regimes to consider the thermal cloud density as static, as assumption already implicitly made (through the use of a fluctuation-dissipation relation) in writing the full stochastic equation in the above reduced form. In that case, a typical simulation would first obtain the self-consistent static thermal equilibrium of a dipolar condensate (e.g. by extending the Hartree-Fock self-consistent method mentioned earlier, Eqs.(32)-(34), to dipolar interactions), and then presumably assume that, under weak perturbations of the thermal cloud, such density remains approximately static [^11] In the context of the contact-interaction ZNG model, this is referred to as the ‘static thermal cloud’ approximation [166]; note, however, that within such an approximation various dynamical dissipations, e.g. vortex decay, were found to have been underestimated in the case of gases with local interactions. This could potentially lead to a useful approximate regime under some weakly-perturbed experimental conditions, although one needs to be mindful that the strong coupling induced by the non-local interactions may still lead to non-negligible indirect thermal cloud dynamics via their dipolar-induced interactions with the non-equilibrium condensate.
@@
$i\frac{\partial\Phi_{0}(x)}{\partial t}=\left(1-i\gamma\right)$ $\left[-\frac{1}{2m}\nabla^{2}+V_{ext}(x)+\frac{C_{dd}}{4\pi}\int d^{3}{\bf r^{\prime}}(n\cdot\nabla)^{2}\frac{1}{|{\bf r}-{\bf r^{\prime}}|}|\Phi_{0}({\bf r^{\prime}})|^{2}\right.$ (85)
@@
Such an equation corresponds precisely to the equation used in [193] to study finite-temperature dipolar condensates, with an ad hoc (constant) value for $\gamma$ (selected in such work as $\gamma=0.1$, which is considerably larger than typical analytically-estimated values). Note that inclusion of quasiparticle physics was achieved by populating the initial state in the simulations by randomised Bogoliubov particle population.
As mentioned earlier, the growing interest in dipolar droplets, where the quantum LHY corrections play a dominant part in the system properties, has led to a modified form of the above equation, with the explicit (but somewhat ad hoc within the premises of such stochastic models) addition of the previously mentioned $\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}$ term on the RHS of Eqs. (85)-(85), leading to the following extended stochastic GPE:
$i\frac{\partial\Phi_{0}(x)}{\partial t}=\left(1-i\gamma\right)$ $\left[-\frac{1}{2m}\nabla^{2}+V_{ext}(x)+\frac{C_{dd}}{4\pi}\int d^{3}{\bf r^{\prime}}(n\cdot\nabla)^{2}\frac{1}{|{\bf r}-{\bf r^{\prime}}|}|\Phi_{0}({\bf r^{\prime}})|^{2}\right.$
@@
Firstly, we note here that our detailed analysis, when all approximations leading to Eqs. (85) - (85) are performed, does in fact give an explicit ab initio expression for the value of $\gamma$ [Eq. (83)], but such an expression needs to be numerically self-consistently evaluated in the context of a particular thermal cloud density approximation, such as the Rayleigh-Jeans distribution in the Hartree-Fock limit. Secondly, it is also pertinent to remind the reader here that the analytical values of $\gamma$ calculated are in fact most likely only reasonable up to an order of magnitude level, rather than an exact, estimate – despite some interesting evidence [118, 131, 104] in the context of weakly-interacting alkali gases that these analytical expressions for an effective $\gamma$ parameter can even lead to quantitative agreement with experiments. Nonetheless, as in most weakly-interacting gas cases, a more direct route for the choice of the effective parameter $\gamma$ is to choose it such that some dynamical observable is consistent with experimental observations – and then assume that such value of $\gamma$ remains fixed and produces other system dynamics fairly accurately, so that the equation acquires further predictive powers. This is precisely the approach used in [192] and is, realistically, the best that one can currently do numerically without explicitly solving the S(P)GPE coupled to the quantum Boltzmann equation.
@@
Secondly, and more importantly, all the earlier comments about the self-consistency of first deriving an effective set of equations when thermal fluctuations dominate, and then supplementing it with corrections known to be dominating the quantum-dominated limit, still hold. While we have no reason to argue against trying such an approach, which is indeed the first logical step in that direction in the absence of more concrete numerically-available models, it is important to highlight that there could be competing assumptions/approximations within any formal derivation including both thermal and quantum effects.
For example, we note in passing here that – at least based on our approach – a term of the LHY form $\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}$ can only arise self-consistently within our formalism when including further fluctuations in the slow felds of higher order in $\Phi^{q}$ and $V_{d}^{q}$. Any such terms may – in addition to the LHY contributions – also generate further additional terms in our studied equations of motion, which might also need to be considered on an equal footing as the LHY correction, unless credible arguments to the contrary could be formulated to discard such extra contributions (which is of course also plausible). This is of course a very challenging question, and remains an open topic of research. Our aim here is not to discourage exploration of such effective stochastic equations – which can evidently provide useful insight into dipolar condensates and a range of experimental observations [192] – but to caution the reader when interpreting results, particularly novel and unexpected ones, emerging from such simulations.
V.4.1 Consistency of the LHY Correction Term

### v2

@@ V.3.1 Stochastic Approach of Stoof
Before commenting further on the use of such an equation in Sec. V.3.3, we note here that an equation like (45) was also obtained as the limiting case of an alternative approach, briefly discussed below.
V.3.2 Stochastic Approach of Gardiner et al.
@@
$iR=-\frac{\beta}{4}\Sigma_{(c)}^{K}(\varepsilon_{q}-\mu)\hskip 28.45274pt{\rm and}\hskip 28.45274pt\Pi^{R}(x,\mathbf{k})-\Pi^{A}(x,\mathbf{k})=\frac{\beta}{2}\Pi^{K}(x,\mathbf{k})\varepsilon_{k}\;.$ (49)
@@
In the SPGPE formalism, the contribution associated with $G(r)$ is known as a growth term (consistent with Stoof’s earlier discussion), and corresponds to an energy- and momentum-conserving binary collision which leads to a change in the number of particles within the $\Phi_{0}$ band. The second term, associated with $M$, corresponds to a collision between a coherent band and an incoherent band particle, which shifts the energies, but does not lead to net population transfer between the two bands. As such, these terms – whose explicit expressions can be found in [80] – have been respectively referred to as ‘number-damping’ and ‘energy-damping’ contributions [160, 161, 162, 163]. Their relative importance is expected to depend on the particular physical scenario being considered, e.g. energy damping (and its associated noise) typically dominate for near-equilibrium systems, except at low phase-space density [246].
@@
it is easy to see that both formalisms become equivalent, upon also dropping the explicit appearance of the projector of the SPGPE formalism. Note that as both ourselves and Stoof are working in a path integral formalism, such projectors do not appear explicitly, but the existence of these two distinct subspaces is implicit from the outset within the path integral approach, a feature becoming relevant in numerical implementations [^15] For example, explicit inclusion of a projector sets a natural band limit to the noise and is important in any such numerical implementations, which will be addressed in future work. Note that the damping parameters depend on the cutoff, which also feeds into the reservoir parameters, with the cutoff usually set at around $2-3\mu$..
Ignoring the so-called scattering term effectively reproduces Eq. (45), but now with the projector, $\mathcal{P}$, appearing explicitly in the equation; further redefining $\psi\to\psi e^{i\mu t}$ and $dW_{G}\to dW_{G}\,e^{i\mu t}$, thus leads to the so-called ‘simple growth’ SPGPE in the form
@@
V.3.3 Stochastic (Projected) Gross-Pitaevskii Predictions
@@
V.3.4 The Dissipative (or Phenomenologically Damped) Gross-Pitaevskii Equation
@@
VI Corresponding Finite-Temperature Equations for Dipolar Gases
In the previous section, we have explicitly shown how our generalised stochastic formalism provides an appealing, more transparent, unifying way, to obtain a broad range of established theories for weakly-interacting atomic gases with contact interactions as limiting cases, based on different approximations. Most notably, our formalism can be seen as creating a unifying framework simultaneously incorporating both stochastic growth and scattering processes on the one hand, and quantum Boltzmann thermal particle dynamics on the other hand, in a dynamically self-consistent manner. In this section we demonstrate the full power of our methodology, by extending our model to systems interacting with long-range dipolar interactions.
VI.1 The Gross-Pitaevskii Equation
@@
[…]
VI.1.1 The Extended Gross-Pitaevskii Equation (eGPE) for Dipolar Condensates
Following here directly the work of Lima and Pelster [23, 24]) (see Appendix C for details), we linearize the condensate wavefunction, following the usual Bogoliubov-de Gennes analysis, via
@@ VI.1.1 The Extended Gross-Pitaevskii Equation (eGPE) for Dipolar Condensates
This section has outlined an established proof of how the equation for the pure condensate mean-field can be systematically extended to account for the role of quantum fluctuations, in the regime where such terms become important. However, it is pertinent to remind the reader that all other formal considerations within this work have focussed on the rather distinct regime where thermal fluctuations dominate. It is therefore plausible to expect that such two distinct pictures, valid in separate limits of the relative ratio of importance of thermal to quantum effects, can be somehow combined, and we comment on this further below (Sec. VI.3).
VI.2 Gross-Pitaevskii-Boltzmann, or ZNG Model of Dipolar Condensates
@@
We stress that the equations presented in this section arise directly from a formalism explicitly focussing on thermal fluctuations: in the regime where such thermal fluctuations are important, leading to corrections on the otherwise pure-condensate physics, it might be reasonable to expect that quantum fluctuations will typically be sub-dominant. As such, it might be that the most relevant regimes in ultracold dipolar systems are either those with condensate and thermal fluctuations (such as the dipolar ZNG equations of Eqs. (71)-(73)), or those with condensate and quantum fluctuations (such as the extended GPE of Eq. (68)), essentially choosing between the generalised finite-temperature models self-consistently derived in an ab initio manner in our work, or, at lower temperatures, to the established LHY quantum correction term summarized in Sec. VI.1.1 and leading to the extended GPE of Eq. (68).
@@
VI.3 An Extended Gross-Pitaevskii, Collisionless Boltzmann Equation Model?
Considering the somewhat ‘orthogonal’ directions of travel presented in Sec. VI.1.1 (quantum fluctuations) and Sec. VI.2 (thermal fluctuations) – each expected to be valid in the relevant regime –, and noting that the latter approach on which our main equations are based does not include condensate fluctuations to all orders, one might consider postulating a combination of those two pictures, albeit in a somewhat ad hoc manner, to interpolate between such regimes.
@@ VI.3 An Extended Gross-Pitaevskii, Collisionless Boltzmann Equation Model?
[…]
Alternatively, observation of Eq. (77) shows that, for $C_{dd}/3\ll g$, or equivalently $\epsilon_{dd}\ll 1$, and a reasonable intermediate temperature $0<T<T_{c}$, the quantum fluctuation term may only pose a minor correction to the overall system properties and dynamics, as expected deeply within the superfluid regime.
The interesting and pertinent question of course remains, whether the above postulated combination of quantum and thermal contributions [Eqs. (74)-(76)] (see also related work in Refs. [47, 50]), which is known to work well in the respective quantum-dominated and thermal-dominated limits may actually provide valid predictions even in the intermediate regime when ${\cal F}\sim O(1)$. In the absence of any other advanced theoretical models addressing this, it may thus be relevant to investigate that numerically, which we defer to a future publication.
VI.4 (Simple Growth) Stochastic Gross-Pitaevskii Equation for Dipolar Condensates
@@
As discussed earlier [Sec. V] for the weakly-interacting gas limit, such an equation is not useful to work with on its own, as the term $iR$ is dynamical, encompassing the dynamics of the incoherent modes.
@@
where the dissipation parameter $\gamma$ is defined here as
@@
One important point that emerges from this equation relates to the explicit appearance of the non-condensate density contribution $\tilde{n}$ in Eqs. (83)-(83). Previously, in the contact-interaction setting, we argued that while any quantitative comparison to experiments strictly requires that contribution to be included, ignoring , to lowest approximation, such a contribution might still give reasonable (but perhaps only qualitative) results, with such treatment possibly also able to model experiments (semi)quantitatively when further supplemented with an appropriately fit $\gamma$ parameter, obtained from experimental measurements. Contrary to that, in the dipolar case we feel it is pertinent to highlight that the situation is more complicated. The reason for this is that the effective dipolar potential $V_{d}$, included in its integrated form in Eq. (83), actually depends critically on the thermal density itself, due to the long-range nature of the interactions. As such, it is not a priori obvious that there exist interesting finite-temperature regimes where ignoring such contributions can be valid. A natural question of course arises how such non-condensate densities could in fact be included in the treatment. Borrowing ideas from earlier treatments, it might be valid in some regimes to consider the thermal cloud density as static, as assumption already implicitly made (through the use of a fluctuation-dissipation relation) in writing the full stochastic equation in the above reduced form. In that case, a typical simulation could, for example, first obtain the self-consistent static thermal equilibrium of a dipolar condensate (e.g. by an extended dipolar Hartree-Fock-like self-consistent method), and then presumably assume that, under weak perturbations of the thermal cloud, such density remains approximately static [^18] In the context of the contact-interaction ZNG model, this is referred to as the ‘static thermal cloud’ approximation [235]; note, however, that within such an approximation various dynamical dissipations, e.g. vortex decay, were found to have been underestimated in the case of gases with local interactions. This could potentially lead to a useful approximate regime under some weakly-perturbed experimental conditions, although one needs to be mindful that the strong coupling induced by the non-local interactions may still lead to non-negligible indirect thermal cloud dynamics via their dipolar-induced interactions with the non-equilibrium condensate.
@@
$i\frac{\partial\Phi_{0}(x)}{\partial t}=\left(1-i\gamma\right)$ $\hskip-8.5359pt\left[-\frac{1}{2m}\nabla^{2}+V_{ext}(x)+\frac{C_{dd}}{4\pi}\int d^{3}{\bf r^{\prime}}(n\cdot\nabla)^{2}\frac{1}{|{\bf r}-{\bf r^{\prime}}|}|\Phi_{0}({\bf r^{\prime}})|^{2}\right.$ (86)
@@
Such an equation corresponds precisely to the equation used in [54] to study finite-temperature dipolar condensates, with an ad hoc (constant) value for $\gamma$ (selected in such work as $\gamma=0.1$, which is considerably larger than typical analytically-estimated values). Note that, in such work, inclusion of quasiparticle physics was achieved by populating the initial state in the simulations by randomised Bogoliubov particle population.
As mentioned earlier, the growing interest in dipolar droplets, where the quantum LHY corrections appear to play a dominant part in the system properties, has led to a modified form of the above equation, with the explicit (but somewhat ad hoc within the premises of such stochastic models) addition of the previously mentioned $\gamma_{QF}|\Phi_{0}|^{3}\Phi_{0}$ term on the RHS of Eqs. (86)-(86), leading to the following extended stochastic GPE:
$i\frac{\partial\Phi_{0}(x)}{\partial t}=\left(1-i\gamma\right)\hskip-1.42271pt$ $\left[-\frac{1}{2m}\nabla^{2}+V_{ext}(x)+\frac{C_{dd}}{4\pi}\int d^{3}{\bf r^{\prime}}(n\cdot\nabla)^{2}\frac{1}{|{\bf r}-{\bf r^{\prime}}|}|\Phi_{0}({\bf r^{\prime}})|^{2}\right.$
@@
Firstly, we note here that our detailed analysis, when all approximations leading to Eqs. (86) - (86) are performed, does in fact give an explicit ab initio expression for the value of $\gamma$ [Eq. (84)], but such an expression needs to be numerically self-consistently evaluated in the context of a particular thermal cloud density approximation, such as the Rayleigh-Jeans distribution in the Hartree-Fock limit. Secondly, it is also pertinent to remind the reader here that the analytical values of $\gamma$ calculated are in fact most likely only reasonable up to an order of magnitude level, rather than an exact, estimate – despite some interesting evidence [178, 191, 163] in the context of weakly-interacting alkali gases that these analytical expressions for an effective $\gamma$ parameter can even lead to quantitative agreement with experiments. Nonetheless, as in most weakly-interacting gas cases, a more direct route for the choice of the effective parameter $\gamma$ is to choose it such that some dynamical observable is consistent with experimental observations – and then assume that such value of $\gamma$ remains fixed and produces other system dynamics fairly accurately, so that the equation acquires further predictive powers. This is precisely the approach used in [57] and is possibly the most extended treatment one can currently do numerically to model the $T>0$ dipolar gas dynamics, without explicitly solving the S(P)GPE coupled to the quantum Boltzmann equation.
@@
Secondly, and more importantly, all the earlier comments about the self-consistency of first deriving an effective set of equations when thermal fluctuations dominate, and then supplementing it with corrections known to be dominating the quantum-dominated limit, still hold. While we have no reason to argue against trying such an approach, which is indeed the first logical step in that direction in the absence of more concrete numerically-available models, it is important to highlight that there could be competing assumptions/approximations within any formal derivation including both thermal and quantum effects, which could further modify such equation.
This is of course a very challenging question, and remains an open topic of research. While acknowledging that such effective stochastic equations can evidently provide useful insight into dipolar condensate experimental observations [57], we nonetheless caution the reader when interpreting novel (or unexpected) results emerging from such simulations.
VI.4.1 Consistency of the LHY Correction Term